# Universality of gravitational radiation from magnetar magnetospheres

### v1

@@ Abstract
The intense magnetic fields inferred from magnetars suggest they may be strong gravitational-wave emitters. Although emissions due to hydromagnetic deformations are more promising from a detection standpoint, exterior fields also contribute a strain. However, numerical evidence suggests that the free energy of stable magnetospheric solutions cannot exceed a few tens of percent relative to the potential state, implying that the gravitational-wave luminosity cannot differ significantly between models. This prompts ‘universality’, in the sense that the strain provides a direct probe of the near-surface field without being muddied by magnetospheric currents. Using a suite of three-dimensional, force-free, general-relativistic solutions for dipole and dipole-plus-quadrupole fields, we find that space-based interferometers may enable marginal detections out to $\lesssim$ kpc distances for slowly-rotating magnetars with fields of $\gtrsim 10^{15}$ G.
@@ I Introduction
Another channel for GW emission comes from the external field directly [13, 14, 15, 16]. Although weaker than many predictions for internally-sourced contributions, the absence of fluid degrees of freedom means that these signals would carry different information. Importantly, magnetospheric solutions are typically classified by their twist relative to the potential (i.e. current-free) state. These currents, sourced via the internal-to-external transfer of helicity and crustal-platelet slippage, are often invoked to explain magnetar activity. For example, hotspot phenomenology may be tied to toroidal field evolution [17, 18], emission spectra and Alfvén wave transport are mediated by current bundles [19, 20], and radio activity is likely regulated by near-surface fluxes [21, 22] as are overstraining ‘failures’ in the outer crust [23, 24].
Despite the diversity of phenomena described above we show, by building a sequence of three-dimensional (3D) and general-relativistic (GR) twisted magnetospheres following Stefanou et al. [25] (henceforth SSP25), that the external GW signal is universal. Because there is a maximum excess energy, relative to an untwisted state, that the magnetosphere can sustain before the (elliptic) force-free problem no longer admits (stable) solutions [26, 27, 28], we argue that strains cannot vary much between models. The magnitude of magnetospheric GW emissions is therefore set mainly by the near-surface field and stellar radius in some combination. A future detection could thus provide a probe of these quantities, which are especially difficult to infer in high-B objects due to their intrinsically noisy nature [29, 22, 30].
@@ II Static magnetospheres
×$\bm{\nabla}\times(e^{\nu}\bm{B})=\alpha\bm{B},$ (1)
@@
meaning that it should be constant along field lines. In the above and throughout, we work with Schwarzschild coordinates such that =$\bm{\nabla}=\{e^{-\lambda}{\partial_{r}},r^{-1}{\partial_{\theta}},+r^{-1}\csc\theta{\partial_{\phi}}\}$ denotes the 3-gradient operator associated with the line element
@@ II.1 Numerical methods
An example of field lines produced using PINNs for $\alpha_{0}=1.5R^{-1},\theta_{1}=\pi/4$, and $\sigma=0.2$ is shown in Figure 1 near the surface (top panel) and at large radii (bottom) for a canonical compactness $\mathcal{C}=0.17$. Key features of the twist are clear: there is a toroidal pressure that inflates the magnetosphere in the right hemisphere relative to the left. The maximum value of $\alpha=1.5R^{-1}$ is reached at the base of the ‘spots’ defined by $\theta_{1}$, with relative cross-sectional width controlled by $\sigma$. These parameters correspond roughly to a pair of spots of size $\sim 2$ km, typical for magnetars (see figure 12 in Ref. [37]). Note that the bottom-panel axes extend to $r\approx 15R$, highlighting the extent of toroidal blow-out.
@@ II.2 Maximum twists
Aside from non-existence, multiple solutions for given boundary data can exist when $\alpha(\bm{x})$ is not constant, with some branches admitting total energies which can reach double that of the potential state (see, for instance, figure 8 in Ref. [43]). Kojima [44] found even larger excesses in axisymmetric, GR magnetospheres for very compact stars (factor $\sim$ few for $\mathcal{C}=0.25$). These ultra-energised solutions are characterised by the presence of disconnected magnetic domains or ‘islands’, seen in studies where the field is twisted via star-disk interactions [45, 46], accretion-induced magnetic burial [47, 48], or crustal currents from magnetothermal evolutions [49]. However, numerical evidence indicates that only the lowest-energy branch is stable, with plasmoids being ejected by kink or buoyant instabilities [50, 20, 28]. Even if large excesses can exist mathematically, they are unlikely to be realised in Nature therefore: the free energy is limited to be of order $\approx 30\%$ in cases of axisymmetric twists – which are always ‘global’ in some sense – or $\approx 5\%$ in realistic systems with localised twists [51, 25]. The inclusion of rotation does not seem to alleviate the problem as magnetic surfaces tend to spread radially unless arrested by hydrodynamic pressures a lá ‘magnetic towers’ [45]. Moreover, SSP25 found that the maximum energy excess of the stable branch is roughly constant as a function of compactness.
@@ III Gravitational radiation from magnetospheric fields
The characteristic GW luminosity of a source is typically quantified in terms of its gravitational multipole moments. While several studies have investigated the polarisation properties of magnetospheric signals [14, 15, 16], we are here primarily interested in the amplitude itself for which knowledge of the strain, $h(t)$, is sufficient. The best-suited detector for Galactic magnetars with spin frequencies $\sim 1$ Hz is DECIGO [31]. The design specifications of this space-based instrument indicates that the sensitivity bowl peaks at such frequencies. Detectability results are deferred to Sec. V.
@@ III.1 Wave strains and quadrupole moments
for the leading-order ($\ell=m=2$) mass quadrupole moment $Q^{22}$, (linear) spin frequency $f$, source distance $d$, Newton’s constant $G$, and speed of light $c$. The first of these quantities is typically that which is the most uncertain in astrophysical applications. In general, $Q^{22}$ has various contributions that appear at different post-Newtonian (PN) orders involving integrals over the components of the Landau-Lifshitz pseudotensor [53]. At Newtonian order, only the $00$-component of the $\mathcal{O}(c^{0})$ piece of the stress-energy enters (i.e., the mass density), which typically dominates for the interior problem. Note that in realistic applications where $h_{0}$ is tracked, one can instead consider functions of time for the relevant variables, i.e., $f(t)$ and possibly $Q^{22}(t)$.
@@
×$\times\left[\cos^{2}\phi\left(1-\cos^{2}\theta\right)-\cos^{2}\theta\right]\Bigg{\}},$
@@
$\approx-2.2\times 10^{38}B_{15}^{2}R_{6}^{5}\left(1+0.28M_{1.4}R_{6}^{-1}\right)\text{ g cm}^{2},$
where the first approximation assumes $R_{\rm LC}/R\gg 1$ – as expected for a slow magnetar – and we have introduced the (CGS units) normalisations $B_{15}=B_{0}/(10^{15}\text{ G})$, $M_{1.4}=M/1.4M_{\odot}$, and $R_{6}=R/(10^{6}\text{ cm})$. Note that the interior fluid deformation of a magnetar with a dipolar magnetic field is of order $Q^{22}_{\rm fluid}\sim 10^{40}B_{15}^{2}M_{1.4}^{-2}R_{6}^{4}\text{ g cm}^{2}$ for poloidal dipoles (see, e.g., equation 17 in Ref. [2]). The hydromagnetic value may thus be larger than that of the magnetosphere (10), as expected. However, if there is a superconducting state such that Meissner expulsion has occurred – often stipulated to match magnetothermal models with X-ray luminosities – the field in the core is effectively zero and $Q^{22}_{\rm fluid}$ reduces by a volumetric factor. In such a case, expression (10) may actually dominate [13, 12]. Modest toroidal fields also reduce $Q^{22}_{\rm fluid}$ as the prolate distortion competes with the oblate one induced by the poloidal field. Curiously, we see that the magnetospheric $Q^{22}$ is negative for poloidal fields, reflecting the fact that $B^{2}$ scales with negative powers of radius rather than positive ones in the interior (as demanded by regularity towards the origin).
@@ IV Results for twisted models
Fixing $\theta_{1}=\pi/4$ and $\sigma=0.2$, quadrupole moments for sequences of varying $\alpha_{0}$ and $\mathcal{C}$ are shown in Figure 4, which can be compared with the left-hand panel of Fig. 3. Graphical representations for some of these twist profiles are shown in figure 4 in SSP25. Again we see that, even for extreme twists $(\alpha_{0}\sim 2.5R^{-1})$ where the force-free approximation is borderline violated, $Q^{22}$ is smaller than expression (10) by at most $\approx 4\%$. For ultra-compact stars with $\mathcal{C}=0.25$ the departure is greatest, while for Newtonian stars we see an uptick because toroidal pressures become more dilute (cf. figures 4d–f in SSP25). The smallness of these changes demonstrate that, regardless of compactness or twist, we expect universality in the sense that the magnetospheric quadrupole moment is roughly constant for fixed values of $B_{0}^{2}R^{5}$ and $\mathcal{C}$.
Although we have not presented an exhaustive sweep, the conclusions are robust: the maximum free energy that a stable magnetosphere can hold is at most of a few tens of percent independently of the stellar compactness. This excess applies to axisymmetric cases, where we find $Q^{22}_{0}/Q^{22}\approx 1.2$ at the limit of the stable branch (see figure 3 in SSP25). For the arguably more realistic cases we have focussed on here with non-axisymmetric, localised spots, the energy discrepancy is at most of order $\approx 5\%$. In general, increases in $Q^{22}$ correlate with the energy excess; this can be seen by comparing Fig. 3 with figure 5 in SSP25 showing relative energies. For example, the $\alpha_{0}=2.5R^{-1}$ case has an excess of $\sim 3\%$ and the quadrupole moment decreases by approximately the same factor. We may therefore anticipate, at most, a factor $\lesssim 1.2$ enhancement for a (likely unrealistic) case that is in some sense maximally twisted with respect to all parameter choices (cf. figure 3 in SSP25). This indicates that GWs emitted by twisted magnetospheres are practically indistinguishable from untwisted ones, and an eventual detection could therefore be used to directly probe the combination $B_{0}^{2}R^{5}$ appearing in expression (10).
@@ IV.1 Dipole-plus-quadrupole fields
where we remind the reader that $\kappa$ is defined as the relative strength between the magnetic quadrupole and dipole components. Note that (11) grows indefinitely with increasing $\kappa$ because the field strength itself grows with $\kappa$. If instead we fix the surface field-strength maximum to a value $\tilde{B}$ (say), then we have $B_{0}\approx\tilde{B}/(1+\kappa)$ depending on the compactness and thus, for quadrupole-dominated systems, $Q^{22}_{0,\kappa\to\infty}/Q^{22}_{0,0}\approx 0.3$. The intermediate regime is instead shown in Figure 5, indicating a local minimum around $\kappa\sim 3$ and that for $\kappa>0$ we have a decrease in the mass-quadrupole moment for fixed surface-maximum $\tilde{B}$. The absolute difference as a function of $\kappa$ is smaller as $\mathcal{C}$ increases because $B^{2}$ becomes more concentrated near the surface (see expression 5) and, having fixed $\tilde{B}$, larger radii contribute less to the integral. The overall value of $Q^{22}_{0,\kappa}$ is lower for more compact stars for the same reason.

### v2

@@ Abstract
The intense magnetic fields inferred from magnetars suggest they may be strong gravitational-wave emitters. Although emissions due to hydromagnetic deformations are more promising from a detection standpoint, exterior fields also contribute a strain. However, numerical evidence suggests that the free energy of stable magnetospheric solutions cannot exceed a few tens of percent relative to the potential state, implying that the magnetospheric contribution to the gravitational-wave luminosity cannot differ significantly between models. This prompts ‘universality’, in the sense that the strain provides a direct probe of the near-surface field without being muddied by magnetospheric currents. Using a suite of three-dimensional, force-free, general-relativistic solutions for dipole and dipole-plus-quadrupole fields, we find that space-based interferometers may enable marginal detections out to $\lesssim$ kpc distances for slowly-rotating magnetars with fields of $\gtrsim 10^{15}$ G independently of internal deformations.
@@ I Introduction
Another channel for GW emission comes from the external field directly [13, 14, 15, 16]. Although weaker than many predictions for internally-sourced contributions, the absence of fluid degrees of freedom means that these signals would carry different information and effectively set a floor value to the GW luminosity. Importantly, magnetospheric solutions are typically classified by their twist relative to the potential (i.e. current-free) state. These currents, sourced via the internal-to-external transfer of helicity and crustal-platelet slippage, are often invoked to explain magnetar activity. For example, hotspot phenomenology may be tied to toroidal field evolution [17, 18], emission spectra and Alfvén wave transport are mediated by current bundles [19, 20], and radio activity is likely regulated by near-surface fluxes [21, 22] as are overstraining ‘failures’ in the outer crust [23, 24].
Despite the diversity of phenomena described above we show, by building a sequence of three-dimensional (3D) and general-relativistic (GR) twisted magnetospheres following Stefanou et al. [25] (henceforth SSP25), that the external GW signal is universal. Because there is a maximum excess energy, relative to an untwisted state, that the magnetosphere can sustain before the (elliptic) force-free problem no longer admits (stable) solutions [26, 27, 28], we argue that strains cannot vary much between models. The magnitude of magnetospheric GW emissions is therefore set mainly by the near-surface field and stellar radius in some combination, while the polarisation phase portrait is set by the distribution of surface currents defining the boundary conditions [15, 14, 16]. A future detection could thus provide a probe of these quantities, which are especially difficult to infer in high-B objects due to their intrinsically noisy nature [29, 22, 30]. We focus on magnetospheric signals in this paper, ignoring the fluid deformation (which may be larger).
@@ II Static magnetospheres
$\bm{\nabla}\times(e^{\nu}\bm{B})=\alpha\bm{B},$ (1)
@@
meaning that it should be constant along field lines. In the above and throughout, we work with Schwarzschild coordinates such that $\bm{\nabla}=\{e^{-\lambda}{\partial_{r}},r^{-1}{\partial_{\theta}},+r^{-1}\csc\theta{\partial_{\phi}}\}$ denotes the 3-gradient operator associated with the line element
@@ II.1 Numerical methods
An example of field lines produced using PINNs for $\alpha_{0}=1.5R^{-1},\theta_{1}=\pi/4$, and $\sigma=0.2$ is shown in Figure 1 near the surface (top panel) and at large radii (bottom) for a canonical compactness $\mathcal{C}=0.17$. Key features of the twist are clear: there is a toroidal pressure that inflates the magnetosphere in the right hemisphere relative to the left. The maximum value of $\alpha=1.5R^{-1}$ is reached at the base of the ‘spots’ defined by $\theta_{1}$, with relative cross-sectional width controlled by $\sigma$. These parameters correspond roughly to a pair of spots of size $\sim 2$ km, typical for magnetars (see figure 12 in Beloborodov and Li [37]). Note that the bottom-panel axes extend to $r\approx 15R$, highlighting the extent of toroidal blow-out.
@@ II.2 Maximum twists
Aside from non-existence, multiple solutions for given boundary data can exist when $\alpha(\bm{x})$ is not constant, with some branches admitting total energies which can reach double that of the potential state (see, for instance, figure 8 in Ref. [43]). Kojima [44] found even larger excesses in axisymmetric, GR magnetospheres for very compact stars (factor $\sim$ few for $\mathcal{C}=0.25$). These ultra-energised solutions are characterised by the presence of disconnected magnetic domains or ‘islands’, seen in studies where the field is twisted via star-disk interactions [45, 46], accretion-induced magnetic burial [47, 48], or crustal currents from magnetothermal evolutions [49]. However, numerical evidence indicates that only the lowest-energy branch is stable (see Appendix C in SSP25), with plasmoids being ejected by kink or buoyant instabilities [50, 20, 28]. Even if large excesses can exist mathematically, they are unlikely to be realised in Nature therefore: the free energy is limited to be of order $\approx 30\%$ in cases of axisymmetric twists – which are always ‘global’ in some sense – or $\approx 5\%$ in realistic systems with localised twists [51, 25]. The inclusion of rotation does not seem to alleviate the problem as magnetic surfaces tend to spread radially unless arrested by hydrodynamic pressures a lá ‘magnetic towers’ [45]. Moreover, SSP25 found that the maximum energy excess of the stable branch is roughly constant as a function of compactness.
@@ III Gravitational radiation from magnetospheric fields
The characteristic GW luminosity of a source is typically quantified in terms of its gravitational multipole moments. While several studies have investigated the polarisation properties of magnetospheric signals [14, 15, 16], we are here primarily interested in the amplitude itself for which knowledge of the strain, $h(t)$, is sufficient (though see also Sec. III.2). The best-suited detector for Galactic magnetars with spin frequencies $\lesssim 1$ Hz is DECIGO [31]. The design specifications of this space-based instrument indicates that the sensitivity bowl peaks at such frequencies. Detectability results are deferred to Sec. V.
@@ III.1 Wave strains and quadrupole moments
for the leading-order ($\ell=m=2$) mass quadrupole moment $Q^{22}$, (linear) spin frequency $f$, source distance $d$, Newton’s constant $G$, and speed of light $c$. The first of these quantities is typically that which is the most uncertain in astrophysical applications. In general, $Q^{22}$ has various contributions that appear at different post-Newtonian (PN) orders involving integrals over the components of the Landau-Lifshitz pseudotensor [53]. At Newtonian order, only the $00$-component of the $\mathcal{O}(c^{0})$ piece of the stress-energy enters (i.e., the mass density), which is often thought to dominate in the interior [54, 55]. Note that in realistic applications where $h_{0}$ is tracked, one can instead consider functions of time for the relevant variables, i.e., $f(t)$ and possibly $Q^{22}(t)$.
@@
$\times\left[\cos^{2}\phi\left(1-\cos^{2}\theta\right)-\cos^{2}\theta\right]\Bigg\},$
@@
[…]
where the first approximation assumes $R_{\rm LC}/R\gg 1$ – as expected for a slow magnetar – and we have introduced the (CGS units) normalisations $B_{15}=B_{0}/(10^{15}\text{ G})$, $M_{1.4}=M/1.4M_{\odot}$, and $R_{6}=R/(10^{6}\text{ cm})$. Curiously, we see that the magnetospheric $Q^{22}$ is negative for poloidal fields, reflecting the fact that $B^{2}$ scales with negative powers of radius rather than positive ones in the interior (as demanded by regularity towards the origin).
@@
III.2 Comparison with hydromagnetic deformations
Many studies of equilibria in the interior of neutron stars stipulate a larger quadrupole moment than that of the magnetosphere (10). The characteristic strain associated with the interior follows that of expression (7) except with the fluid quadrupole moment [52],
$h_{0}^{\rm fluid}=\frac{16\pi^{2}G}{c^{4}}\frac{|Q^{22}_{\rm fluid}|f^{2}}{d}.$ (11)
Notably, X-ray [54, 55] and radio [58] modulations have been interpreted as evidence for precession and large quadrupole moments in some magnetars. For example, high-cadence radio monitoring of the magnetar XTE J1810–197 following its outburst in 2018 led Desvignes et al. [58] to infer a quadrupole moment of order $|Q^{22}_{\rm fluid}|\gtrsim 10^{38}\text{ g cm}^{2}$. Given the (GR-corrected) polar field strength estimate of this object of $B_{p}\sim 10^{14}$ G [25], we may anticipate a ratio $|Q^{22}_{\rm magneto}|/|Q^{22}_{\rm fluid}|\lesssim 10^{-2}$ from expression (10).
Even if the magnetospheric contribution is comparatively weak however, its detectability is largely independent of the interior. This is because its polarisation pattern (i.e., the phase portraits of $h_{+}$ and $h_{\times}$) is set by the distribution of surface currents [14, 15, 16], while that of the interior is determined by the equation of state, poloidal-toroidal partition, and other factors [4, 52]. For example, in the simple case of a biaxial deformation and ignoring detector-frame corrections, the plus polarisation profile of the total system reads [14]
$h_{+}\approx$ $h^{\rm magneto}_{0}\sin\vartheta_{0}\Big[\sin\vartheta_{0}\left(1+\cos^{2}\iota\right)\cos\left(2\Omega\tau+2\varphi_{0}\right)$ (12)
$+\cos\vartheta_{0}\cos\iota\sin\iota\sin\left(\Omega\tau+\varphi_{0}\right)\Big]$
$+\frac{1}{2}h^{\rm fluid}_{0}\left(1+\cos^{2}\iota\right)\cos\left(2\Omega\tau\right),$
where $\iota$ is the inclination from the rotation axis to the line-of-sight, $\Omega=2\pi/P$ is the angular frequency, $\tau$ is the retarded time, and $\vartheta_{0}$ and $\varphi_{0}$ respectively denote the angles between the Cartesian $z$- and $x$-axes with the magnetic axis in the rotating frame. An important feature evident in expression (12) is that the phase offset $\varphi_{0}$ is present only in the 1PN waveform, $h_{0}^{\rm magneto}(t)$. As explained by Nazari and Roshan [14], it arises because the PN phase portrait is set by gradients of the magnetic pressure alone, while the hydromagnetic term depends on the total Lorentz force. The distinction becomes even more apparent for tangled fields in the interior, as explored in detail by Lasky and Melatos [4], where the star may also radiate at other harmonics of $\Omega\tau$ depending on the Euler angles.
Over the course of a multi-year observational campaign with DECIGO, the superposition of terms within (12) and $h_{\times}(t)$ could therefore be independently visible in the full waveform if $\varphi_{0}\neq 0$. As the detectability of $h_{0}^{\rm fluid}$ has been explored extensively in the literature (see, e.g., Refs. [1, 5, 6, 4, 3, 2, 7, 8, 9, 10, 11, 12]), we henceforth ignore it and consider only the detectability of the (independent) component $h_{0}^{\rm magneto}$ defined by expression (7). Although we focus on the amplitude (7) in this paper to provide sky- and polarisation-averaged estimates (Sec. V), the above should be kept in mind since, even if the hydromagnetic deformation dominates, the magnetospheric signal will not be subsumed in general. In fact, a detection of both $h_{0}^{\rm fluid}(t)$ and $h_{0}^{\rm magneto}(t)$ could be used to independently constrain the poloidal-toroidal partition and near-surface field strength, which would provide more information than the detection of an individual piece [59].
@@ IV Results for twisted models
Fixing $\theta_{1}=\pi/4$ and $\sigma=0.2$, quadrupole moments for sequences of varying $\alpha_{0}$ and $\mathcal{C}$ are shown in Figure 4, which can be compared with the left-hand panel of Fig. 3. Graphical representations for some of these twist profiles are shown in figure 4 in SSP25. Again we see that, even for extreme twists $(\alpha_{0}\sim 2.5R^{-1})$ where the force-free approximation is borderline violated, $Q^{22}$ is smaller than expression (10) by at most $\approx 3\%$. For ultra-compact stars with $\mathcal{C}=0.25$ the departure is greatest, while for Newtonian stars we see an uptick because toroidal pressures become more dilute (cf. figures 4d–f in SSP25). The smallness of these changes demonstrate that, regardless of compactness or twist, we expect universality in the sense that the magnetospheric quadrupole moment is roughly constant for fixed values of $B_{0}^{2}R^{5}$ and $\mathcal{C}$.
Although we have not presented an exhaustive sweep, the conclusions are robust: the maximum free energy that a stable magnetosphere can hold is at most of a few tens of percent independently of the stellar compactness. This excess applies to axisymmetric cases, where we find $Q^{22}_{0}/Q^{22}\approx 1.2$ at the limit of the stable branch (see figure 3 in SSP25). For the arguably more realistic cases we have focussed on here with non-axisymmetric, localised spots, the energy discrepancy is at most of order $\approx 5\%$. In general, increases in $Q^{22}$ correlate with the energy excess, as is clear from expression (8). For example, the $\alpha_{0}=2.5R^{-1}$ case has an excess of $\sim 3\%$ and the quadrupole moment decreases by approximately the same factor. We may therefore anticipate, at most, a factor $\lesssim 1.2$ enhancement for a (likely unrealistic) case that is in some sense maximally twisted with respect to all parameter choices (cf. figure 3 in SSP25). This indicates that GWs emitted by twisted magnetospheres are practically indistinguishable from untwisted ones, and an eventual detection could therefore be used to directly probe the combination $B_{0}^{2}R^{5}$ appearing in expression (10).
@@ IV.1 Dipole-plus-quadrupole fields
where we remind the reader that $\kappa$ is defined as the relative strength between the magnetic quadrupole and dipole components. Note that expression (13) grows indefinitely with increasing $\kappa$ because the field strength itself grows with $\kappa$. If instead we fix the surface field-strength maximum to a value $\tilde{B}$ (say), then we have $B_{0}\approx\tilde{B}/(1+\kappa)$ depending on the compactness and thus, for quadrupole-dominated systems, $Q^{22}_{0,\kappa\to\infty}/Q^{22}_{0,0}\approx 0.3$. The intermediate regime is instead shown in Figure 5, indicating a local minimum around $\kappa\sim 3$ and that for $\kappa>0$ we have a decrease in the mass-quadrupole moment for fixed surface-maximum $\tilde{B}$. The absolute difference as a function of $\kappa$ is smaller as $\mathcal{C}$ increases because $B^{2}$ becomes more concentrated near the surface (see expression 5) and, having fixed $\tilde{B}$, larger radii contribute less to the integral. The overall value of $Q^{22}_{0,\kappa}$ is lower for more compact stars for the same reason.